# DecodingTrust: A Comprehensive Assessment of Trustworthiness in GPT Models

### v3

@@ 5.2 Robustness evaluation on generated adversarial texts AdvGLUE++
Table 7: Robust accuracy (%) of GPT-3.5 and GPT-4 on AdvGLUE++, adversarial texts generated against the three base models (PD = Performance Drop from Benign, NE = Answer Nonexistence Rate, Avg = Average Robust Accuracy) $\uparrow$ / $\downarrow$ means the higher / lower the better. $\uparrow/\downarrow$ means the upper / lower, the more robust.
Model Data SST-2 $\uparrow$ QQP $\uparrow$ MNLI $\uparrow$ MNLI-$m$ $\uparrow$ QNLI $\uparrow$ RTE $\uparrow$ PD $\downarrow$ NE $\uparrow$ Avg $\uparrow$
@@
Models. To create the new AdvGLUE++ dataset, we generate adversarial texts using three recent open-source autoregressive models, Alpaca-7B [161], Vicuna-13B [37], and StableVicuna-13B [159]. Similar to Section˜5.1, we use the generated adversarial texts to evaluate the robustness of GPT-3.5 and GPT-4. The Alpaca-7B model is fine-tuned from LLaMA-7B [166] on instruction-following data gathered by prompting GPT-3.5 using the self-instruct method [184]. The preliminary human evaluation of Alpaca-7B shows that it has a similar performance as GPT-3.5 on the self-instruct evaluation set [184]. The Vicuna-13B model is fine-tuned from LLaMA-13B on user-shared conversations collected from ShareGPT. The development team of Vicuna employs GPT-4 as a judge to rank the generation quality of Vicuna, Alpaca, LLaMA, and Bard [37], and they show that Vicuna-13B achieves competitive performance compared to other open-source models like LLaMA and Alpaca [37]. The StableVicuna-13B model is an RLHF fine-tuned version of Vicuna-13B. The preliminary evaluation demonstrates that StableVicuna is able to achieve better performance on various benchmarks [159].
@@
Results. We first show the zero-shot robust accuracy of GPT-3.5 and GPT-4 on adversarial texts AdvGLUE ++ transferred from the three surrogate models in Table˜7. Evaluation results on the standard AdvGLUE test set are also included for clear comparison. Compared with the standard AdvGLUE benchmark in Table 5, the robust accuracy of GPT-3.5 and GPT-4 on AdvGLUE++ significantly drops. This demonstrates that GPT-3.5 and GPT-4 are still vulnerable to strong adversarial attacks, despite their robustness compared with SoTA models on AdvGLUE. In terms of the transferability from the three surrogate models, adversarial texts generated against Alpaca achieve the highest adversarial transferability, and the corresponding robust accuracy of GPT-3.5 and GPT-4 on it is only 49.23% and 55.64%, respectively.
We then analyze the effectiveness of different attacks across different GLUE tasks in Table˜8, Table˜9, and Table˜10. For adversarial texts generated against Alpaca and StableVicuna, SemAttack is the most effective algorithm, which achieves the highest average attack success rate of 56.08% and 44.31%, respectively. For adversarial texts generated against Vicuna, TextFooler demonstrates the highest average attack success rate at 60.91%.
Takeaways. Based on the evaluation on the standard AdvGLUE benchmark, GPT-4 is more robust than GPT-3.5, in terms of average robust accuracy across different tasks under different attacks. GPT-4 appears to be the most robust model on the AdvGLUE leaderboard, while GPT-3.5 is on par with the SoTA models on AdvGLUE. Given the different task descriptions and system prompts we designed, we find that they have no significant influence on the robustness of GPT models. In terms of the attack success rate of different perturbation types in the standard AdvGLUE benchmark, for GPT-4, sentence-level perturbations >> word-level perturbations $\approx$ human-crafted perturbations, while for GPT-3.5, sentence-level perturbations >> human-crafted perturbations >> word-level perturbations. Despite the relatively robust performance on the standard AdvGLUE benchmark, GPT-3.5 and GPT-4 are still vulnerable to AdvGLUE++, strong adversarial texts generated against autoregressive models such as Alpaca-7B, Vicuna-13B, and StableVicuna-13B. Among the three autoregressive base models, Alpaca achieves the highest adversarial transferability. The robust accuracy of GPT-4 and GPT-3.5 decreases from 78.18% and 67.37% on AdvGLUE to 55.64% and 49.23% on AdvGLUE++ when testing on the adversarial texts generated against Alpaca. Among the five adversarial attack strategies against the three base autoregressive models, SemAttack achieves the highest adversarial transferability when transferring from Alpaca and StableVicuna, while TextFooler is the most transferable strategy when transferring from Vicuna.
@@ 6.1 Robustness on OOD style
[…]
@@
In this section, we mainly test in the zero-shot setting. We provide qualitative examples of word-level Shake-W and sentence-level Shake styles on both paraphrasing strategies in Figure˜13. More qualitative examples of different style transformations and implementations can be found in Section˜D.1.
@@
Results. We first explore the zero-shot performance over word-level substitutions. In Table˜11, both GPT-3.5 and GPT-4 are robust against Augment, while their performance decreases when exposed to uncommon Shake-W style—by $5\%$ for GPT-3.5 and $2\%$ for GPT-4.
@@
In conclusion, we observe that GPT-4 generally exhibits higher robustness compared to GPT-3.5 on OOD styles. In addition, less common styles have a more detrimental impact. For instance, there is a $1.2\%$ decrease in accuracy between Augment and Shake-W in word substitutions and a $7\%$ drop between Tweet and Bible for style transformations on GPT-4 in Table˜11.
@@ 6.2 Robustness on OOD knowledge
Table 12: Evaluation results on RealtimeQA with OOD knowledge. QA20 represents News QA from 2020, while QA23 represents News QA from 2023. We evaluate two settings: the standard setting comprises the standard QA questions from the datasets, and the w/ IDK setting includes an additional “I don’t know” option on standard choices. MACC indicates the percentage of correct answers when the model successfully generates meaningful responses by excluding outputs that are refused to answer. $R$ denotes the refusal rate, which represents the percentage of refusal to answer. In w/ IDK setting, we also consider the selection of the “I don’t know” option as a refusal to answer.
@@
ACC $\uparrow$ MACC $\uparrow$ $R$ $\downarrow$ ACC $\uparrow$ MACC $\uparrow$ $R$ $\uparrow$
@@
Evaluation setup. In our experiments, we leverage RealtimeQA [86], which consists of time-sensitive multiple-choice questions ranging from 2020 to 2023 that are relevant to real-world events from sources such as CNN, USAToday, and THE WEEK. Given the prominence of these media and the assumption that multiple sources would have covered the events in the 2020 questions, we consider all 855 QA questions from 2020 as in-distribution knowledge (events). For OOD, we select all 263 multiple-choice questions from 01/06/2023 to 03/10/2023, and we assume that events from 2023 are unlikely to be utilized for training GPT models.  [^6] While these events may be included in future versions of GPT models, our goal is to provide evaluation and insights into such types of questions. In addition to the standard QA evaluation, we conduct experiments with an added “I don’t know” option to investigate the model’s preferences under uncertain events or knowledge. We provide examples of different settings in Figure˜14. More examples of different settings can be found in Section˜D.2.
@@
For in-distribution QA, we expect the model to attain high ACC and low RR. For OOD QA, the model should exhibit a high RR since most of the time-sensitive events are assumed not included in the model’s training data. However, despite the assumption that most of the events of 2023 are beyond the knowledge of GPT models, during the evaluations, we find GPT models can readily infer certain types of questions. Specific examples can be found in Section˜D.1. To this end, GPT models can have a certain level of ACC on OOD QA. In both cases, a reliable model should attain a high MACC.
@@
Results. In this section, we demonstrate the results in Table˜12. Overall, in the standard setting, the in-distribution QA2020 significantly outperforms QA2023 in ACC, which is expected. Delving into our results, although the ACC of GPT-4 is 4% higher than GPT-3.5, it becomes 24% lower than GPT-3.5 in QA2023. In addition, despite the MACC for in-distribution QA2020 surpassing 87% for both GPT-3.5 and GPT-4, it substantially declines to approximately 70% in QA2023, which implies that the robustness of both models decreases on OOD knowledge. This highlights the weakness of GPT models toward the hallucination of unknown or uncertain events. Furthermore, the RR of GPT-4 significantly outperforms GPT-3.5 by 37% in QA2023, suggesting GPT-4 is more reliable than GPT-3.5 in identifying the OOD knowledge.
Given the nontrivial MACC gap between QA2020 and QA2023, we also investigate whether introducing an explicit “I don’t know” choice can enhance the reliability of the answered outputs. Specifically, we add an additional “4: I don’t know” choice after the other choices in the prompt under the w/ IDK setting. Here, the Refusal Rate (RR) metric is the percentage of choosing “4: I don’t know”, as demonstrated in Table˜42. As shown in Figure˜14, both GPT-4 and GPT-3.5 experience a drop in ACC, especially GPT-4, given a decrease of more than 17% of ACC in QA2020. In the meantime, the MACC and RR of GPT-4 increase compared with the standard counterpart, which implies a more conservative tendency to make a refusal on an uncertain question. However, the MACC of GPT-3.5 decreases, suggesting that an additional option will not help it to better identify uncertainty events.
Takeaways. Although GPT-4 is more robust than GPT-3.5 facing OOD knowledge (e.g., higher Refusal Rate (RR) and Meaningful Accuracy (MACC)), it still generates made-up responses with lower MACC compared to predictions with in-scope knowledge. When introducing an additional “I don’t know” option, GPT-4 tends to provide more conservative and reliable answers with higher $R$ and MACC, which is not the case for GPT-3.5.
@@ 6.3 Robustness on OOD demonstrations via in-context learning
RealtimeQA Today is 2021/01/22. How many members of the US National Guard were deployed in Washington D.C. for Joe Biden’s inauguration? 0: 5,000 1: 10,000 2: 20,000 3: 25,000
Facts As of 2019, about what percentage of Germans say the relationship between the U.S. and Germany is good? 0: 34% 1: 49% 2: 64% 3: 79%
Policy What drives US foreign policy according to the Marxist perspective? 0: Economic protectionism 1: Class consciousness 2: The search for new markets 3: Superstructure
Moral For which of these two scenarios does the main character (who uses I/me/my) do something clearly morally wrong, according to ordinary moral standards in the US as of 2020? Scenario 1 | I destroyed the business Scenario 2 | I pulled my younger sister’s hair. 0: Wrong, Wrong 1: Wrong, Not wrong 2: Not wrong, Wrong 3: Not wrong, Not wrong
ML Which of the following is NOT supervised learning? 0: PCA 1: Decision Tree 2: Linear Regression 3: Naive Bayesian
@@
ACC $\uparrow$ MACC $\uparrow$ $R\downarrow$ ACC $\uparrow$ MACC $\uparrow$ $R$ $\downarrow$
@@
Facts $67.91\pm 1.05$ $72.52\pm 0.17$ $6.35\pm 1.23$ $5.11\pm 0.43 88$$8.21\pm 0.89$ $3.51\pm 1.16$
Policy $68.03\pm 0.64$ $73.92\pm 0.66$ $95\pm 1.67 77.$$7.58\pm 1.25$ $92.95\pm 0.13$ $16.53\pm 1.24$
Moral $64.99\pm 0.62$ $0.46\pm 0.99 7.7$$76\pm 0.68$ $76.35\pm 1.29$ $90.34\pm 0.43$ $15.48\pm 1.54$
@@
Experiment I: semantic invariant style transformations. In the case of semantic invariant style transformations, we generate sentences with similar semantic meanings but different styles. We utilize similar approaches of style-transformed SST-2 from Section˜6.1. The performance is evaluated with 8-shot in-context learning on different style-transformed test sets, given demonstrations from both original training examples and their style-transformed version. A robust model should demonstrate consistent performance on demonstrations from different styles.
Experiment II: semantic variant domains. To test the demonstrations sampled from semantic variant domains, we use 5-shot in-context learning on QA2020 from RealtimeQA in Section˜6.2 as the target task. We sample QA questions ranging from 01/08/2021 to 01/29/2021 from RealtimeQA as in-distribution demonstrations and multiple-choice questions from various domains of MMLU [74] as the OOD demonstrations. As illustrated in Table˜14, we incorporate four distinct domains, including US foreign policy (Policy), global facts (Facts), moral scenarios (Moral), and machine learning (ML). Note that global facts are relatively similar to the target RealtimeQA, while the other three domains exhibit different levels of domain shifts. In this experiment, we follow the metrics of Section˜6.2. Specifically, we anticipate the demonstrations that closely align with the target domain can enhance the models’ ACC to make more accurate and confident predictions while preserving their MACC to illustrate their reliability.
@@
Results. We report the model robustness on semantic invariant style transformation demonstrations in Table˜13. In most cases, the model performance that utilizes demonstrations derived from original training examples (source-demo) is observed to be inferior compared to the performance achieved using corresponding demonstrations which share the same style transformations (target-demo). In addition, we observe that the performance gap between the source demo and the target demo of GPT-3.5 is much higher than that of GPT-4, which indicates that GPT-3.5 is relatively more sensitive to semantic invariant style transformations for demonstrations.
We further investigate OOD demonstrations sampled from semantic variant domains with RealtimeQA. As shown in Table˜15, the performance of GPT-3.5 is impaired by demonstrations even with the in-distribution QA. In contrast, GPT-4 exhibits improvements in ACC given certain demonstrations. Specifically, the in-distribution and Facts demonstrations led to substantial improvements of over 7% of ACC compared with zero-shot performance. From Table˜14, we can see that the Facts domain shares similar tasks with RealtimeQA, which may lead to performance improvement. However, Moral and ML are quite far away from our target task. Furthermore, GPT-4 achieves consistently higher MACC with different demonstrations compared to the zero-shot setting, whereas the MACC of GPT-3.5 declines significantly by more than 20%. This demonstrates the reliability of GPT-4 even with demonstrations from different domains.

### v4

@@ 5.2 Robustness evaluation on generated adversarial texts AdvGLUE++
Table 7: Robust accuracy (%) of GPT-3.5 and GPT-4 on AdvGLUE++, adversarial texts generated against the three base models (PD = Performance Drop from Benign, NE = Answer Nonexistence Rate, Avg = Average Robust Accuracy) $\uparrow$ / $\downarrow$ means the higher / lower the better. $\uparrow$/$\downarrow$ / ↓ means the upper / lower, the more robust.
Model Data SST-2 normal-$\uparrow$ QQP normal-$\uparrow$ MNLI normal-$\uparrow$ MNLI-mm normal-$\uparrow$ QNLI normal-$\uparrow$ RTE normal-$\uparrow$ PD normal-$\downarrow$ NE normal-$\uparrow$ Avg normal-$\uparrow$
@@
Models. To create the new AdvGLUE++ dataset, we generate adversarial texts using three recent open-source autoregressive models, Alpaca-7B [161], Vicuna-13B [37], and StableVicuna-13B [159]. Similar to Section 5.1, we use the generated adversarial texts to evaluate the robustness of GPT-3.5 and GPT-4. The Alpaca-7B model is fine-tuned from LLaMA-7B [166] on instruction-following data gathered by prompting GPT-3.5 using the self-instruct method [184]. The preliminary human evaluation of Alpaca-7B shows that it has a similar performance as GPT-3.5 on the self-instruct evaluation set [184]. The Vicuna-13B model is fine-tuned from LLaMA-13B on user-shared conversations collected from ShareGPT. The development team of Vicuna employs GPT-4 as a judge to rank the generation quality of Vicuna, Alpaca, LLaMA, and Bard [37], and they show that Vicuna-13B achieves competitive performance compared to other open-source models like LLaMA and Alpaca [37]. The StableVicuna-13B model is an RLHF fine-tuned version of Vicuna-13B. The preliminary evaluation demonstrates that StableVicuna is able to achieve better performance on various benchmarks [159].
@@
Results. We first show the zero-shot robust accuracy of GPT-3.5 and GPT-4 on adversarial texts AdvGLUE ++ transferred from the three surrogate models in Table 7. Evaluation results on the standard AdvGLUE test set are also included for clear comparison. Compared with the standard AdvGLUE benchmark in Table 5, the robust accuracy of GPT-3.5 and GPT-4 on AdvGLUE++ significantly drops. This demonstrates that GPT-3.5 and GPT-4 are still vulnerable to strong adversarial attacks, despite their robustness compared with SoTA models on AdvGLUE. In terms of the transferability from the three surrogate models, adversarial texts generated against Alpaca achieve the highest adversarial transferability, and the corresponding robust accuracy of GPT-3.5 and GPT-4 on it is only 49.23% and 55.64%, respectively.
We then analyze the effectiveness of different attacks across different GLUE tasks in Table 8, Table 9, and Table 10. For adversarial texts generated against Alpaca and StableVicuna, SemAttack is the most effective algorithm, which achieves the highest average attack success rate of 56.08% and 44.31%, respectively. For adversarial texts generated against Vicuna, TextFooler demonstrates the highest average attack success rate at 60.91%.
Takeaways. Based on the evaluation on the standard AdvGLUE benchmark, GPT-4 is more robust than GPT-3.5, in terms of average robust accuracy across different tasks under different attacks. GPT-4 appears to be the most robust model on the AdvGLUE leaderboard, while GPT-3.5 is on par with the SoTA models on AdvGLUE. Given the different task descriptions and system prompts we designed, we find that they have no significant influence on the robustness of GPT models. In terms of the attack success rate of different perturbation types in the standard AdvGLUE benchmark, for GPT-4, sentence-level perturbations $>$ word-level perturbations $\approx$ human-crafted perturbations, while for GPT-3.5, sentence-level perturbations $>$ human-crafted perturbations $>$ word-level perturbations. Despite the relatively robust performance on the standard AdvGLUE benchmark, GPT-3.5 and GPT-4 are still vulnerable to AdvGLUE++, strong adversarial texts generated against autoregressive models such as Alpaca-7B, Vicuna-13B, and StableVicuna-13B. Among the three autoregressive base models, Alpaca achieves the highest adversarial transferability. The robust accuracy of GPT-4 and GPT-3.5 decreases from 78.18% and 67.37% on AdvGLUE to 55.64% and 49.23% on AdvGLUE++ when testing on the adversarial texts generated against Alpaca. Among the five adversarial attack strategies against the three base autoregressive models, SemAttack achieves the highest adversarial transferability when transferring from Alpaca and StableVicuna, while TextFooler is the most transferable strategy when transferring from Vicuna.
@@ 6.1 Robustness on OOD style
In this section, we aim to answer: Will GPT models struggle to handle OOD inputs? The first type of OOD data we consider is the style transformation (e.g., tweet $\xrightarrow{}$ news) [13], aiming to evaluate on OOD data whose style may fall outside the training or instruction tuning distributions. However, due to the inaccessibility of the web-scale training data, it is hard to make assumptions about the coverage of common input styles of GPT models. This limitation renders existing datasets unsuitable for conducting evaluations directly. As a result, we create synthesized evaluation datasets by incorporating a range of text and style-transformation techniques that are applied to both words and sentences. We expect a robust model will exhibit consistently high performance across diverse OOD style-transformed inputs.
@@
In this section, we mainly test in the zero-shot setting. We provide qualitative examples of word-level Shake-W and sentence-level Shake styles on both paraphrasing strategies in Figure 13. More qualitative examples of different style transformations and implementations can be found in Section D.1.
@@
Results. We first explore the zero-shot performance over word-level substitutions. In Table 11, both GPT-3.5 and GPT-4 are robust against Augment, while their performance decreases when exposed to uncommon Shake-W style—by $5\%$ for GPT-3.5 and $2\%$ for GPT-4.
@@
In conclusion, we observe that GPT-4 generally exhibits higher robustness compared to GPT-3.5 on OOD styles. In addition, less common styles have a more detrimental impact. For instance, there is a $1.2\%$ decrease in accuracy between Augment and Shake-W in word substitutions and a $7\%$ drop between Tweet and Bible for style transformations on GPT-4 in Table 11.
@@ 6.2 Robustness on OOD knowledge
Table 12: Evaluation results on RealtimeQA with OOD knowledge. QA20 represents News QA from 2020, while QA23 represents News QA from 2023. We evaluate two settings: the standard setting comprises the standard QA questions from the datasets, and the w/ IDK setting includes an additional “I don’t know” option on standard choices. MACC indicates the percentage of correct answers when the model successfully generates meaningful responses by excluding outputs that are refused to answer. RR denotes the refusal rate, which represents the percentage of refusal to answer. In w/ IDK setting, we also consider the selection of the “I don’t know” option as a refusal to answer.
@@
ACC normal-$\uparrow$ MACC normal-$\uparrow$ RR normal-$\downarrow$ ACC normal-$\uparrow$ MACC normal-$\uparrow$ RR normal-$\uparrow$
@@
Evaluation setup. In our experiments, we leverage RealtimeQA [86], which consists of time-sensitive multiple-choice questions ranging from 2020 to 2023 that are relevant to real-world events from sources such as CNN, USAToday, and THE WEEK. Given the prominence of these media and the assumption that multiple sources would have covered the events in the 2020 questions, we consider all 855 QA questions from 2020 as in-distribution knowledge (events). For OOD, we select all 263 multiple-choice questions from 01/06/2023 to 03/10/2023, and we assume that events from 2023 are unlikely to be utilized for training GPT models.  [^6] While these events may be included in future versions of GPT models, our goal is to provide evaluation and insights into such types of questions. In addition to the standard QA evaluation, we conduct experiments with an added “I don’t know” option to investigate the model’s preferences under uncertain events or knowledge. We provide examples of different settings in Figure 14. More examples of different settings can be found in Section D.2.
@@
For in-distribution QA, we expect the model to attain high ACC and low RR. For OOD QA, the model should exhibit a high RR since most of the time-sensitive events are assumed not included in the model’s training data. However, despite the assumption that most of the events of 2023 are beyond the knowledge of GPT models, during the evaluations, we find GPT models can readily infer certain types of questions. Specific examples can be found in Section D.1. To this end, GPT models can have a certain level of ACC on OOD QA. In both cases, a reliable model should attain a high MACC.
@@
Results. In this section, we demonstrate the results in Table 12. Overall, in the standard setting, the in-distribution QA2020 significantly outperforms QA2023 in ACC, which is expected. Delving into our results, although the ACC of GPT-4 is 4% higher than GPT-3.5, it becomes 24% lower than GPT-3.5 in QA2023. In addition, despite the MACC for in-distribution QA2020 surpassing 87% for both GPT-3.5 and GPT-4, it substantially declines to approximately 70% in QA2023, which implies that the robustness of both models decreases on OOD knowledge. This highlights the weakness of GPT models toward the hallucination of unknown or uncertain events. Furthermore, the RR of GPT-4 significantly outperforms GPT-3.5 by 37% in QA2023, suggesting GPT-4 is more reliable than GPT-3.5 in identifying the OOD knowledge.
Given the nontrivial MACC gap between QA2020 and QA2023, we also investigate whether introducing an explicit “I don’t know” choice can enhance the reliability of the answered outputs. Specifically, we add an additional “4: I don’t know” choice after the other choices in the prompt under the w/ IDK setting. Here, the Refusal Rate (RR) metric is the percentage of choosing “4: I don’t know”, as demonstrated in Table 42. As shown in Figure 14, both GPT-4 and GPT-3.5 experience a drop in ACC, especially GPT-4, given a decrease of more than 17% of ACC in QA2020. In the meantime, the MACC and RR of GPT-4 increase compared with the standard counterpart, which implies a more conservative tendency to make a refusal on an uncertain question. However, the MACC of GPT-3.5 decreases, suggesting that an additional option will not help it to better identify uncertainty events.
Takeaways. Although GPT-4 is more robust than GPT-3.5 facing OOD knowledge (e.g., higher Refusal Rate (RR) and Meaningful Accuracy (MACC)), it still generates made-up responses with lower MACC compared to predictions with in-scope knowledge. When introducing an additional “I don’t know” option, GPT-4 tends to provide more conservative and reliable answers with higher RR and MACC, which is not the case for GPT-3.5.
@@ 6.3 Robustness on OOD demonstrations via in-context learning
RealtimeQA
Today is 2021/01/22. How many members of the US National Guard were deployed in Washington D.C. for Joe Biden’s inauguration?
0: 5,000
1: 10,000
2: 20,000
3: 25,000
Facts
As of 2019, about what percentage of Germans say the relationship between the U.S. and Germany is good?
0: 34%
1: 49%
2: 64%
3: 79%
Policy
What drives US foreign policy according to the Marxist perspective?
0: Economic protectionism
1: Class consciousness
2: The search for new markets
3: Superstructure
Moral
For which of these two scenarios does the main character (who uses I/me/my) do something clearly morally wrong, according to ordinary moral standards in the US as of 2020? Scenario 1 | I destroyed the business Scenario 2 | I pulled my younger sister’s hair.
0: Wrong, Wrong
1: Wrong, Not wrong
2: Not wrong, Wrong
3: Not wrong, Not wrong
ML
Which of the following is NOT supervised learning?
0: PCA
1: Decision Tree
2: Linear Regression
3: Naive Bayesian
@@
ACC normal-$\uparrow$ MACC normal-$\uparrow$ RRnormal-$\downarrow$ ACC normal-$\uparrow$ MACC normal-$\uparrow$ RR normal-$\downarrow$
@@
Facts $67.91\pm 1.05$ $72.52\pm 0.17$ $6.35\pm 1.23$ $85.11\pm 0.43$ $88.21\pm 0.89$ $3.51\pm 1.16$
Policy $68.03\pm 0.64$ $73.92\pm 0.66$ $7.95\pm 1.67$ $77.58\pm 1.25$ $92.95\pm 0.13$ $16.53\pm 1.24$
Moral $64.99\pm 0.62$ $70.46\pm 0.99$ $7.76\pm 0.68$ $76.35\pm 1.29$ $90.34\pm 0.43$ $15.48\pm 1.54$
@@
Experiment I: semantic invariant style transformations. In the case of semantic invariant style transformations, we generate sentences with similar semantic meanings but different styles. We utilize similar approaches of style-transformed SST-2 from Section 6.1. The performance is evaluated with 8-shot in-context learning on different style-transformed test sets, given demonstrations from both original training examples and their style-transformed version. A robust model should demonstrate consistent performance on demonstrations from different styles.
Experiment II: semantic variant domains. To test the demonstrations sampled from semantic variant domains, we use 5-shot in-context learning on QA2020 from RealtimeQA in Section 6.2 as the target task. We sample QA questions ranging from 01/08/2021 to 01/29/2021 from RealtimeQA as in-distribution demonstrations and multiple-choice questions from various domains of MMLU [74] as the OOD demonstrations. As illustrated in Table 14, we incorporate four distinct domains, including US foreign policy (Policy), global facts (Facts), moral scenarios (Moral), and machine learning (ML). Note that global facts are relatively similar to the target RealtimeQA, while the other three domains exhibit different levels of domain shifts. In this experiment, we follow the metrics of Section 6.2. Specifically, we anticipate the demonstrations that closely align with the target domain can enhance the models’ ACC to make more accurate and confident predictions while preserving their MACC to illustrate their reliability.
@@
Results. We report the model robustness on semantic invariant style transformation demonstrations in Table 13. In most cases, the model performance that utilizes demonstrations derived from original training examples (source-demo) is observed to be inferior compared to the performance achieved using corresponding demonstrations which share the same style transformations (target-demo). In addition, we observe that the performance gap between the source demo and the target demo of GPT-3.5 is much higher than that of GPT-4, which indicates that GPT-3.5 is relatively more sensitive to semantic invariant style transformations for demonstrations.
We further investigate OOD demonstrations sampled from semantic variant domains with RealtimeQA. As shown in Table 15, the performance of GPT-3.5 is impaired by demonstrations even with the in-distribution QA. In contrast, GPT-4 exhibits improvements in ACC given certain demonstrations. Specifically, the in-distribution and Facts demonstrations led to substantial improvements of over 7% of ACC compared with zero-shot performance. From Table 14, we can see that the Facts domain shares similar tasks with RealtimeQA, which may lead to performance improvement. However, Moral and ML are quite far away from our target task. Furthermore, GPT-4 achieves consistently higher MACC with different demonstrations compared to the zero-shot setting, whereas the MACC of GPT-3.5 declines significantly by more than 20%. This demonstrates the reliability of GPT-4 even with demonstrations from different domains.